OpenStudy (anonymous):

What value of x will make this an arithmetic series? x, 0.5x + 7, 3x - 1...

OpenStudy (inkyvoyd):

HINT: In an arithmetic series, each difference is the same.

OpenStudy (blockcolder):

For that to be an arithmetic series, x must satisfy the ff: (0.5x+7)-x=(3x-1)-(0.5x+7)

OpenStudy (anonymous):

I know that the difference is the same. But I still can't find the value for x.

OpenStudy (anonymous):

If the series is S1,S2,S3,S4,... then, S2-S1 = S3-S2 = S4-S3 = Sn - S(n-1)

OpenStudy (anonymous):

You that means you can set up that equation that blockolder set up for you. Solve that equation for x.

OpenStudy (anonymous):

I still seem to be unable to solve it =S

OpenStudy (inkyvoyd):

You sure?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

x, 0.5x + 7, 3x - 1... Treat it like any normal algebra problem. Simplify. Gather the xs on one side and the constants on the other. (0.5x+7) - x = (3x-1)-(0.5x+7) -0.5x+7 = 2.5x -8 -3x +7 = -8 -3x = - 15 x = (-15/-3) = 5

OpenStudy (anonymous):

what if instead of 3x - 1, we had 3x + 1, what would happen then?